Assume three events A.B.C are such that p(A)=0.5, p(B)=0.6, p(C)=0.4,p(AnB)=0.3,
p(AnC)=0.1, p(BnC)=0.2, and p(AnBnC)=0.05. Find
(1) p(AnBnC')
(2) p(AuBuC)
(3) p(A'nBnC')

Found 3 solutions by Edwin McCravy, ikleyn, greenestamps:
Answer by Edwin McCravy(20077)

Assume three events A.B.C are such that p(A)=0.5, p(B)=0.6, p(C)=0.4,p(AnB)=0.3,
p(AnC)=0.1, p(BnC)=0.2, and p(AnBnC)=0.05. Find
(1) p(AnBnC')
(2) p(AuBuC)
(3) p(A'nBnC')
We draw a Venn diagram with 3 sets ( the 3 circles) and a universal set (the big rectangle).  
We let the letters d through k indicate the probabilities of the 8 regions:

   

Since the probability of something in the Universal set occurring 
is 1, referring to the above Venn diagram:

p(U) = 1.0 = d+e+f+g+h+i+j+k
p(A) = 0.5 = d+e+g+h 
p(B) = 0.6 = e+f+h+i 
p(C) = 0.4 = g+h+i+j
p(AnB) = 0.3 = e+h 
p(AnC) = 0.1 = g+h
p(BnC) = 0.2 = f+i
p(AnBnC) = 0.05 = h

So we have this system of equations:
(1)     d+e+f+g+h+i+j+k = 1.0
(2)     d+e  +g+h       = 0.5 
(3)       e+f  +h+i     = 0.6 
(4)           g+h+i+j   = 0.4
(5)       e    +h       = 0.3 
(6)           g+h       = 0.1
(7)             h+i     = 0.2
(8)             h       = 0.05 

Use (8) to substitute 0.05 for h in (7) to get i = 0.15,
and in (6) to get g = 0.05, and in (5) to get e = 0.25.

Substitute for g,h and i in (4) to get j = 0.15
Substitute for e,h and i in (3) to get f = 0.15
Substitute for e,g and h in (2) to get d = 0.15
Substitute for d,e,f,g,h,i,j in (1) to get k = 0.05

So we have:

    

(1) p(AnBnC') = e = 0.25
(2) p(AuBuC) = h = 0.05  <--that was given!!!
(3) p(A'nBnC') = f = 0.15

Edwin

Answer by ikleyn(53614)

.

            It is MUCH easier and simpler solution than that by Edwin.

            It requires you to apply only basic formulas of the elementary set theory.


(1)   P(A n B n C')

To calculate  P(A n B n C'), notice that  (A n B n C')  are those elements  (events)  of (A n B) that do not belong to C;

in other words,  (A n B n C') = (A n B) \ (A n B n C),      (1)

where the sign " \ " denotes subtraction of sets (of subsets).


Then equality (1) implies that

    P(A n B n C') = P(A n B) - P(A n B n C) = (substitute the given data) = 0.3 - 0.05 = 0.25.        ANSWER


(2)   P(A U B U C)

Use the formula

    P(A U B U C) = P(A) + P(B) + P(C) - P(AnB) - P(AnC) - P(BnC) + P(A n B n C)     (2)

which is valid for any subsets  A, B and C  of the universal set U.


    This formula is very well known in elementary set theory.

    If you do not know its proof, here it is in few lines:

    
        Without my explanations, you understand well why I added P(A), P(B) and P(C) in the right side of the formula (2).

        But when I added these terms, I counted the intersections P(AnB), P(AnC)  and P(BnC) twice - so I need to subtract 
        these terms in the right side of the formula (2).

        But again,  when I added the terms P(A), P(B) and P(C) in the right side of the formula (2),
        I counted the intersection P(A n B n C) thrice. When I later subtracted the terms P(AnB), P(AnC)  and P(BnC), I fully compensate it,
        but now this intersection is not covered in the formula at all.
      
        Therefore I must add the term P(A n B n C), and this last step makes everything in equilibrium / (in balance).  

        The proof is completed.



    Now, when the formula (2) is proved, simply substitute the given input data in it. You will get

        P(A U B U C) = 0.5 + 0.6 + 0.4 - 0.3 - 0.1 - 0.2 + 0.05 = 0.95      ANSWER

Notice that the Edwin's answer in this point is incorrect.


(3)   P(A' n B n C')

To calculate  P(A' n B n C'), notice that  (A' n B n C')  are those elements  (events)  of B that do belong NEITHER  A  NOR  C;

in other words,  (A' n B n C') = (B \ (A n B) \ (C n B))    (3)



Then equality (2) implies that

    P(A' n B n C') = P(B \ (A n B) \ (C n B)) = (P(B) - P(AnB)) + (P(B) - P(CnB)) + P(A n B n C)     (4)


Notice that in the right side of the formula (4), I subtracted the probability P(AnB) from P(B), then subtracted P(CnB) from P(B), 
so I subtracted the probability P(AnC) of the intersection (AnC) twice; therefore, I should add P(A n B n C) to restore the equilibrium.


Substitute the given data into the formula (4), you get  the ANSWER

    P(A' n B n C') =   (0.6  - 0.3   - 0.2)   +   0.05       = 0.15.    (5)

----------------

When such problems come to the forum,  a competition starts between Edwin and me.

Edwin's first move is to solve the problem by the most complicated method.  So he sets up the system of linear equations of the order  8
and then solves it.

I,  in opposite,  think that such approach is  IRRELEVANT  and that the problem should be solved by the most simple way,
using only basic notions,  conceptions,  properties and formulas of the elementary set theory.

I am  129%  sure that the problem was designed and intended to be solved exactly in this way.

In any case, doing in this way  (or,  at least reading my solution)  you will learn something useful from the elementary set theory.

----------------

Regarding formula (2), see many other similar solved problems in the lessons
    - Advanced problems on counting elements in sub-sets of a given finite set
    - Challenging problems on counting elements in subsets of a given finite set
in this site.



Answer by greenestamps(13295)



Just a comment on the responses you got from tutors @edwin and @ikleyn....

I certainly would not solve a system of 8 equations in 8 unknowns to solve this problem, as Edwin did... although that might work for some people.

Definitely the much easier way to go is the way ikleyn went.

However, the description of the solution in words, as she showed it, might be hard for you to follow.

If you are new to this kind of problem and aren't yet comfortable with them, I would recommend you look at Edwin's last two diagrams and use them to follow ikleyn's solutions to see that the problem is in fact relatively easy.